OpenStudy (anonymous):

take the derivative of h(x) = e^(x^2 * sin(x))

OpenStudy (anonymous):

is this \[e^{x^2\times \sin(x)}\]?

OpenStudy (anonymous):

if so you get get \[e^{x^2\times \sin(x)}\times (2x\sin(x)+x^2\cos(x))\] by the chain rule and product rule

OpenStudy (anonymous):

Yes it is it's just hard to type it without the equation coding. Thank you!